Find the height to the nearest foot of a tree that makes a 50-foot shadow when a 2 ft. 6 in. pole is used to make a shadow of 3 feet.
Answer by checkley71(8403)

X/50=2.5/3 NOW CROSS MULTIPLY.
3X=2.5*50
3X=125
X=125/3
X=41.667 FEET IS THE HEIGHT OF THE TREE.